OpenStudy (anonymous):

how to solve \sqrt{t - 14} - \sqrt{t+96} = 9

OpenStudy (phi):

t+96 will always be larger than t-14 so it looks like there will be no solution. but we can proceed as follows: square both sides \[ \sqrt{t - 14} - \sqrt{t+96} = 9\\ t-14 + t+96 - 2 \sqrt{(t-14)(t+96)}=81\\ 2t + (96-14-81) = 2 \sqrt{(t-14)(t+96)} \\ 2t+1 = 2 \sqrt{(t-14)(t+96)} \] you can now square both sides again \[ 4t^2 +4t+1 = 4(t-14)(t+96) \\ 4t^2 +4t+1 = 4 ( t^2 +82t - 1344)\\ \cancel{4t^2} +4t+1 = \cancel{4t^2} + 328t - 5376 \\ 5377 = 324 t\\ t= \frac{5377}{324} \approx 16.59567901 \] if we try that solution we get the left hand side = -9 so it is an extraneous solution... and there is no solution